# A contramodule generalization of Neeman’s flat and projective module theorem

### v1

@@ Introduction
Geometrically, flat contramodules over a commutative topological ring with a countable topology base of open ideals can be identified with flat pro-quasi-coherent pro-sheaves on an ind-affine $\aleph_{0}$-ind-scheme [5, Section 7.11.3], [23, Examples 3.8], [26, Example 8.9]. This constitutes a geometric motivation for the present work.
@@ Acknowledgement
The author is supported by the GAČR project 23-05148S and the Czech Academy of Sciences (RVO 67985840).
@@ 1. Accessible Subcategories
Let $\mathsf{K}$ be a category with $\kappa$-directed (equivalently, $\kappa$-filtered) colimits. An object $S\in\mathsf{K}$ is said to be $\kappa$-presentable if the functor $\operatorname{Hom}_{\mathsf{K}}(S,{-})\colon\mathsf{K}\longrightarrow\mathsf{Sets}$ from $\mathsf{K}$ to the category of sets $\mathsf{Sets}$ preserves $\kappa$-directed colimits. When the category $\mathsf{K}$ is (pre)additive, this condition is equivalent to the functor $\operatorname{Hom}_{\mathsf{K}}(S,{-})\colon\mathsf{K}\longrightarrow\mathsf{Ab}$ to the category of abelian groups $\mathsf{Ab}$ preserving $\kappa$-directed colimits.
@@
This well-known result goes back, at least, to [16, Proposition 2.1], [9, Section 4.1], and [15, Proposition 5.11] (notice that what we call finitely accessible categories were called “finitely presented categories” in [9, 15]). The nontrivial part is the “if” implication in the last assertion; all the other assertions follow easily. See [25, Proposition 1.2] for some details. ∎
@@
Part (a) holds because $\kappa$-directed colimits commute with countable limits in the category $\mathsf{Sets}$ or $\mathsf{Ab}$, cf. [26, Lemma 1.1]. Part (b) is a particular case of [25, Theorem 6.2]; cf. [25, proof of Corollary 10.4]. ∎
@@ 2. Contraderived Categories
Let $\mathsf{K}$ be an abelian category, and let $\mathsf{F}$, $\mathsf{C}\subset\mathsf{K}$ be two classes of objects. One denotes by $\mathsf{F}^{\perp_{1}}\subset\mathsf{K}$ the class of all objects $X\in\mathsf{K}$ such that $\operatorname{Ext}^{1}_{\mathsf{K}}(F,X)=0$ for all $F\in\mathsf{F}$. Dually, the notation ${}^{\perp_{1}}\mathsf{C}\subset\mathsf{K}$ stands for the class of all objects $Y\in\mathsf{K}$ such that $\operatorname{Ext}^{1}_{\mathsf{K}}(Y,C)=0$ for all $C\in\mathsf{C}$. Similarly, $\mathsf{F}^{\perp_{\geq 1}}\subset\mathsf{K}$ is the class of all objects $X\in\mathsf{K}$ such that $\operatorname{Ext}^{n}_{\mathsf{K}}(F,X)=0$ for all $n\geq 1$ and $F\in\mathsf{F}$, while ${}^{\perp_{\geq 1}}\mathsf{C}\subset\mathsf{K}$ is the class of all objects $Y\in\mathsf{K}$ such that $\operatorname{Ext}^{n}_{\mathsf{K}}(Y,C)=0$ for all $n\geq 1$ and $C\in\mathsf{C}$.
A full subcategory $\mathsf{F}\subset\mathsf{K}$ is said to be self-generating if for every epimorphism $K\longrightarrow F$ in $\mathsf{K}$ with $F\in\mathsf{F}$ there exists a morphism $G\longrightarrow K$ in $\mathsf{K}$ with $G\in\mathsf{F}$ such that the composition $G\longrightarrow K\longrightarrow F$ is an epimorphism $G\longrightarrow F$ in $\mathsf{K}$. A full subcategory $\mathsf{F}\subset\mathsf{K}$ is said to be generating if for every object $K\in\mathsf{K}$ there exists an object $G\in\mathsf{F}$ together with an epimorphism $G\longrightarrow K$ in $\mathsf{K}$. Clearly, any generating full subcategory is self-generating.
A full subcategory $\mathsf{F}\subset\mathsf{K}$ is said to be self-resolving if it is self-generating and closed under extensions and kernels of epimorphisms. We will say that a full subcategory $\mathsf{F}\subset\mathsf{K}$ is resolving if it is generating and closed under extensions and kernels of epimorphisms. Clearly, any resolving full subcategory is self-resolving.
@@
Let $\mathsf{K}$ be an abelian category and $\mathsf{F}\subset\mathsf{K}$ be a self-generating full subcategory closed under kernels of epimorphisms. Then one has $\mathsf{F}^{\perp_{1}}=\mathsf{F}^{\perp_{\geq 1}}\subset\mathsf{K}$.
@@
This assertion holds more generally for any exact category $\mathsf{K}$ (in the sense of Quillen). The argument from [34, Lemma 6.17] applies. See [3, Lemma 1.3] for some details. ∎
@@
Given a class of objects $\mathsf{S}\subset\mathsf{K}$, we denote by $\mathsf{Fil}(\mathsf{S})\subset\mathsf{K}$ the class of all objects filtered by (objects isomorphic to) the objects from $\mathsf{S}$. A class of objects $\mathsf{F}\subset\mathsf{K}$ is said to be deconstructible if there exists a set of objects $\mathsf{S}\subset\mathsf{K}$ such that $\mathsf{F}=\mathsf{Fil}(\mathsf{S})$. The following result is known classically as the Eklof lemma [10, Lemma 1].
@@
In the stated generality, a proof can be found in [28, Lemma 4.5]; see also [30, Proposition 1.3]. ∎
Given an additive category $\mathsf{B}$, we denote by $\mathsf{Hot}(\mathsf{B})$ the triangulated category of (unbounded) complexes in $\mathsf{B}$ with morphisms up to cochain homotopy. The notation $B^{\text{\smaller\smaller$\scriptstyle\bullet$}}[i]$,  $i\in\mathbb{Z}$ stands for the cohomological degree shift of a complex $B^{\text{\smaller\smaller$\scriptstyle\bullet$}}$; so $(B^{\text{\smaller\smaller$\scriptstyle\bullet$}}[i])^{n}=B^{n+i}$ for all $n\in\mathbb{Z}$. The following lemma is well-known.
@@
See, e. g., [30, Lemma 5.1]. ∎
Let $\mathsf{B}$ be an abelian (or exact) category with enough projective objects. A complex $B^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ in $\mathsf{B}$ is said to be contraacyclic (in the sense of Becker [4, Proposition 1.3.6(1)]) if, for every complex of projective objects $P^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ in $\mathsf{B}$, the complex of abelian groups $\operatorname{Hom}_{\mathsf{B}}(P^{\text{\smaller\smaller$\scriptstyle\bullet$}},B^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ is acyclic. The full subcategory of contraacyclic complexes is denoted by $\mathsf{Ac}^{\mathsf{bctr}}(\mathsf{B})\subset\mathsf{Hot}(\mathsf{B})$ or (depending on context) $\mathsf{Ac}^{\mathsf{bctr}}(\mathsf{B})\subset\mathsf{Com}(\mathsf{B})$. The triangulated Verdier quotient category
@@
Any short exact sequence $0\longrightarrow K^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow L^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow M^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow 0$ in $\mathsf{B}$ can be viewed as a bicomplex with three rows. Then one can consider its total complex, which we will denote by $\operatorname{Tot}(K^{\text{\smaller\smaller$\scriptstyle\bullet$}}\to L^{\text{\smaller\smaller$\scriptstyle\bullet$}}\to M^{\text{\smaller\smaller$\scriptstyle\bullet$}})$.
@@
(a) For any short exact sequence $0\longrightarrow K^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow L^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow M^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow 0$ in $\mathsf{B}$, the total complex $\operatorname{Tot}(K^{\text{\smaller\smaller$\scriptstyle\bullet$}}\to L^{\text{\smaller\smaller$\scriptstyle\bullet$}}\to M^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ is contraacyclic.
@@
These observations seem to go back to [20, Section 3.5]. Part (b) is obvious. In part (a), the point is that, for every complex of projective objects $P^{\text{\smaller\smaller$\scriptstyle\bullet$}}\in\mathsf{B}$, the short seqeuence of complexes of abelian groups $0\longrightarrow\operatorname{Hom}_{\mathsf{B}}(P^{\text{\smaller\smaller$\scriptstyle\bullet$}},K^{\text{\smaller\smaller$\scriptstyle\bullet$}})\longrightarrow\operatorname{Hom}_{\mathsf{B}}(P^{\text{\smaller\smaller$\scriptstyle\bullet$}},L^{\text{\smaller\smaller$\scriptstyle\bullet$}})\longrightarrow\operatorname{Hom}_{\mathsf{B}}(P^{\text{\smaller\smaller$\scriptstyle\bullet$}},M^{\text{\smaller\smaller$\scriptstyle\bullet$}})\longrightarrow 0$ is exact. The total complex of any short exact sequence of complexes of abelian groups is exact. ∎
We refer to the survey paper [7] for general background on exact categories in the sense of Quillen. Given an exact category $\mathsf{E}$, a complex in $\mathsf{E}$ is said to be absolutely acyclic [19, Section 2.1], [20, Section 3.3], [21, Appendix A] if it belongs to the minimal thick subcategory of $\mathsf{Hot}(\mathsf{E})$ containing the total complexes of (admissible) short exact sequences in $\mathsf{E}$. Equivalently, a complex in $\mathsf{E}$ is absolutely acyclic if and only if it belongs to the minimal full subcategory of $\mathsf{Com}(\mathsf{E})$ containing contractible complexes and closed under extensions and direct summands [31, Proposition 8.12].
@@
This is [19, Remark 2.1]. For additional information, see [26, Proposition 6.2]. ∎
Let us denote by $\mathsf{B}_{\mathsf{proj}}\subset\mathsf{B}$ the full subcategory of projective objects in $\mathsf{B}$.
@@
This is [30, Corollary 7.4]. For a generalization, see [31, Corollary 6.14]. ∎

### v3

@@ Introduction
We also consider the exact category of cotorsion $\mathfrak{R}$-contramodules $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}}$. According to [36, Theorem 7.17 and Corollary 7.21], for any associative ring $R$, the inclusion of exact/abelian categories $R{\operatorname{\mathsf{--Mod}}}^{\mathsf{cot}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$ (where $R{\operatorname{\mathsf{--Mod}}}^{\mathsf{cot}}$ is the full subcategory of cotorsion left $R$-modules) induced equivalences of their derived and contraderived categories $\mathsf{D}(R{\operatorname{\mathsf{--Mod}}}^{\mathsf{cot}})\simeq\mathsf{D}(R{\operatorname{\mathsf{--Mod}}})$ and $\mathsf{D}^{\mathsf{bctr}}(R{\operatorname{\mathsf{--Mod}}}^{\mathsf{cot}})\simeq\mathsf{D}^{\mathsf{bctr}}(R{\operatorname{\mathsf{--Mod}}})$. For a complete, separated topological ring $\mathfrak{R}$ with a countable base of neighborhoods of zero consisting of two-sided ideals, it was shown in [34, Corollary 12.8] that the inclusion of exact/abelian categories $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}}\longrightarrow\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ induces an equivalence of the derived categories $\mathsf{D}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}})\simeq\mathsf{D}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$. In this paper we show that the same inclusion induces an equivalence of the contraderived categories $\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}})\simeq\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$.
Geometrically, flat contramodules over a commutative topological ring with a countable topology base of open ideals can be identified with flat pro-quasi-coherent pro-sheaves on an ind-affine $\aleph_{0}$-ind-scheme [5, Section 7.11.3], [31, Examples 3.8], [34, Example 8.9]. This constitutes a geometric motivation for the present work.
This paper consists roughly of two parts. Sections 14 present the preliminary material, and then in Sections 56 we prove contramodule versions of results related to the flat/projective periodicity theorem of Benson–Goodearl and Neeman. Section 7 spells out additional preliminaries for the second part, and then in Sections 810 contramodule versions of results related to the cotorsion periodicity theorem of Bazzoni, Cortés-Izurdiaga, and Estrada are proved. The final Section 11 summarizes the results obtained in Sections 56 and 810.
@@ Acknowledgement
I am grateful to Jan Št’ovíček for helpful discussions. The author is supported by the GAČR project 23-05148S and the Czech Academy of Sciences (RVO 67985840).
@@ 1. Accessible Subcategories
Let $\mathsf{K}$ be a category with $\kappa$-directed (equivalently, $\kappa$-filtered) colimits. An object $S\in\mathsf{K}$ is said to be $\kappa$-presentable if the functor $\operatorname{Hom}_{\mathsf{K}}(S,{-})\colon\mathsf{K}\longrightarrow\mathsf{Sets}$ from $\mathsf{K}$ to the category of sets $\mathsf{Sets}$ preserves $\kappa$-directed colimits. When the category $\mathsf{K}$ is (pre)additive, an object $S\in\mathsf{K}$ is $\kappa$-presentable if and only if the functor $\operatorname{Hom}_{\mathsf{K}}(S,{-})\colon\mathsf{K}\longrightarrow\mathsf{Ab}$ from $\mathsf{K}$ to the category of abelian groups $\mathsf{Ab}$ preserves $\kappa$-directed colimits.
@@
This well-known result goes back, at least, to [24, Proposition 2.1], [11, Section 4.1], and [23, Proposition 5.11] (notice that what we call finitely accessible categories were called “locally finitely presented categories” in [11, 23]). The nontrivial part is the “if” implication in the last assertion; all the other assertions follow easily. See [35, Proposition 1.2] for some details. ∎
@@
Part (a) holds because $\kappa$-directed colimits commute with countable limits in the category $\mathsf{Sets}$ or $\mathsf{Ab}$, cf. [34, Lemma 1.1]. Part (b) is a particular case of [35, Theorem 6.2]; cf. [35, proof of Corollary 10.4]. ∎
@@ 2. Contraderived Categories
We refer to the survey paper [8] for general background on exact categories in the sense of Quillen. The Yoneda Ext groups in an exact category $\mathsf{K}$ can be defined, e. g., as $\operatorname{Ext}^{n}_{\mathsf{K}}(X,Y)=\operatorname{Hom}_{\mathsf{D}(\mathsf{K})}(X,Y[n])$ for all $X$, $Y\in\mathsf{K}$ and $n\geq 0$. Here $\mathsf{D}(\mathsf{K})$ denotes the derived category of $\mathsf{K}$.
Let $\mathsf{K}$ be an exact category, and let $\mathsf{F}$, $\mathsf{C}\subset\mathsf{K}$ be two classes of objects. One denotes by $\mathsf{F}^{\perp_{1}}\subset\mathsf{K}$ the class of all objects $X\in\mathsf{K}$ such that $\operatorname{Ext}^{1}_{\mathsf{K}}(F,X)=0$ for all $F\in\mathsf{F}$. Dually, the notation ${}^{\perp_{1}}\mathsf{C}\subset\mathsf{K}$ stands for the class of all objects $Y\in\mathsf{K}$ such that $\operatorname{Ext}^{1}_{\mathsf{K}}(Y,C)=0$ for all $C\in\mathsf{C}$. Similarly, $\mathsf{F}^{\perp_{\geq 1}}\subset\mathsf{K}$ is the class of all objects $X\in\mathsf{K}$ such that $\operatorname{Ext}^{n}_{\mathsf{K}}(F,X)=0$ for all $n\geq 1$ and $F\in\mathsf{F}$, while ${}^{\perp_{\geq 1}}\mathsf{C}\subset\mathsf{K}$ is the class of all objects $Y\in\mathsf{K}$ such that $\operatorname{Ext}^{n}_{\mathsf{K}}(Y,C)=0$ for all $n\geq 1$ and $C\in\mathsf{C}$.
Given an exact category $\mathsf{K}$, any full subcategory $\mathsf{E}\subset\mathsf{K}$ closed under extensions in $\mathsf{K}$ can be endowed with an exact category structure in which the (admissible) short exact sequences in $\mathsf{E}$ are the short exact sequences in $\mathsf{K}$ with the terms belonging to $\mathsf{E}$. We will say that the exact structure on $\mathsf{E}$ is inherited from the exact structure on $\mathsf{K}$.
A full subcategory $\mathsf{F}\subset\mathsf{K}$ is said to be self-generating if for every admissible epimorphism $K\longrightarrow F$ in $\mathsf{K}$ with $F\in\mathsf{F}$ there exists a morphism $G\longrightarrow K$ in $\mathsf{K}$ with $G\in\mathsf{F}$ such that the composition $G\longrightarrow K\longrightarrow F$ is an admissible epimorphism $G\longrightarrow F$ in $\mathsf{K}$. A full subcategory $\mathsf{F}\subset\mathsf{K}$ is said to be generating if for every object $K\in\mathsf{K}$ there exists an object $G\in\mathsf{F}$ together with an admissible epimorphism $G\longrightarrow K$ in $\mathsf{K}$. Clearly, any generating full subcategory is self-generating.
A full subcategory $\mathsf{F}\subset\mathsf{K}$ is said to be self-resolving if it is self-generating and closed under extensions and kernels of admissible epimorphisms. We will say that a full subcategory $\mathsf{F}\subset\mathsf{K}$ is resolving if it is generating and closed under extensions and kernels of admissible epimorphisms. Clearly, any resolving full subcategory is self-resolving.
@@
Let $\mathsf{K}$ be an exact category and $\mathsf{F}\subset\mathsf{K}$ be a self-generating full subcategory closed under kernels of admissible epimorphisms. Then one has $\mathsf{F}^{\perp_{1}}=\mathsf{F}^{\perp_{\geq 1}}\subset\mathsf{K}$.
@@
The argument from [46, Lemma 6.17] applies. See [3, Lemma 1.3] for some additional details. ∎
@@
Given a class of objects $\mathsf{S}\subset\mathsf{K}$, we denote by $\mathsf{Fil}(\mathsf{S})\subset\mathsf{K}$ the class of all objects filtered by (objects isomorphic to) the objects from $\mathsf{S}$. A class of objects $\mathsf{F}\subset\mathsf{K}$ is said to be deconstructible if there exists a set of objects $\mathsf{S}\subset\mathsf{K}$ such that $\mathsf{F}=\mathsf{Fil}(\mathsf{S})$. The following result is known classically as the Eklof lemma [12, Lemma 1].
@@
In the stated generality, a proof can be found in [38, Lemma 4.5]; see also [40, Proposition 1.3]. ∎
The following lemma is well known and easy.
Lemma 2.3.
Let $\mathsf{B}$ be an abelian category, $A\in\mathsf{B}$ be an object, and $n\in\mathbb{Z}$ be an integer. Denote by $D_{n,n+1}^{\text{\smaller\smaller$\scriptstyle\bullet$}}(A)$ the contractible two-term complex $\dotsb\longrightarrow 0\longrightarrow A\overset{{\mathrm{id}}}{\longrightarrow}A\longrightarrow 0\longrightarrow\dotsb$ concentrated in the cohomological degrees $n$ and $n+1$. Then, for any complex $B^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ in $\mathsf{B}$ and all integers $i\geq 0$ there is a natural isomorphism of abelian groups
$\operatorname{Ext}^{i}_{\mathsf{Com}(\mathsf{B})}(D_{n,n+1}^{\text{\smaller\smaller$\scriptstyle\bullet$}}(A),B^{\text{\smaller\smaller$\scriptstyle\bullet$}})\simeq\operatorname{Ext}^{i}_{\mathsf{B}}(A,B^{n}).$
Proof.
See, e. g., [15, Lemma 3.1(5)] or [33, Lemma 1.8]. ∎
Given an additive category $\mathsf{B}$, we denote by $\mathsf{Hot}(\mathsf{B})$ the triangulated category of (unbounded) complexes in $\mathsf{B}$ with morphisms up to cochain homotopy. The notation $B^{\text{\smaller\smaller$\scriptstyle\bullet$}}[i]$,  $i\in\mathbb{Z}$ stands for the cohomological degree shift of a complex $B^{\text{\smaller\smaller$\scriptstyle\bullet$}}$; so $(B^{\text{\smaller\smaller$\scriptstyle\bullet$}}[i])^{n}=B^{n+i}$ for all $n\in\mathbb{Z}$. The following lemma is well known.
@@
See, e. g., [15, Lemma 2.1], [3, Lemma 1.6] or [33, Lemma 1.9]. ∎
Let $\mathsf{B}$ be an abelian (or exact) category with enough projective objects. A complex $B^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ in $\mathsf{B}$ is said to be contraacyclic (in the sense of Becker [4, Proposition 1.3.6(1)]) if, for every complex of projective objects $P^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ in $\mathsf{B}$, the complex of abelian groups $\operatorname{Hom}_{\mathsf{B}}^{\text{\smaller\smaller$\scriptstyle\bullet$}}(P^{\text{\smaller\smaller$\scriptstyle\bullet$}},B^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ is acyclic. The full subcategory of contraacyclic complexes is denoted by $\mathsf{Ac}^{\mathsf{bctr}}(\mathsf{B})\subset\mathsf{Hot}(\mathsf{B})$ or (depending on context) $\mathsf{Ac}^{\mathsf{bctr}}(\mathsf{B})\subset\mathsf{Com}(\mathsf{B})$. The triangulated Verdier quotient category
@@
We are interested in (termwise admissible) short exact sequences $0\longrightarrow K^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow L^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow M^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow 0$ of complexes in $\mathsf{B}$. Any such short exact sequence can be viewed as a bicomplex with three rows in $\mathsf{B}$. Then one can consider its total complex, which we will denote by $\operatorname{Tot}(K^{\text{\smaller\smaller$\scriptstyle\bullet$}}\to L^{\text{\smaller\smaller$\scriptstyle\bullet$}}\to M^{\text{\smaller\smaller$\scriptstyle\bullet$}})$.
@@
Let $\mathsf{B}$ be an exact category with enough projective objects. Then
(a) for any short exact sequence $0\longrightarrow K^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow L^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow M^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow 0$ of complexes in $\mathsf{B}$, the total complex $\operatorname{Tot}(K^{\text{\smaller\smaller$\scriptstyle\bullet$}}\to L^{\text{\smaller\smaller$\scriptstyle\bullet$}}\to M^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ is contraacyclic;
@@
These observations seem to go back to [28, Section 3.5]. Part (b) is obvious. In part (a), the point is that, for every complex of projective objects $P^{\text{\smaller\smaller$\scriptstyle\bullet$}}\in\mathsf{B}$, the short seqeuence of complexes of abelian groups $0\longrightarrow\operatorname{Hom}_{\mathsf{B}}^{\text{\smaller\smaller$\scriptstyle\bullet$}}(P^{\text{\smaller\smaller$\scriptstyle\bullet$}},K^{\text{\smaller\smaller$\scriptstyle\bullet$}})\longrightarrow\operatorname{Hom}_{\mathsf{B}}^{\text{\smaller\smaller$\scriptstyle\bullet$}}(P^{\text{\smaller\smaller$\scriptstyle\bullet$}},L^{\text{\smaller\smaller$\scriptstyle\bullet$}})\longrightarrow\operatorname{Hom}_{\mathsf{B}}^{\text{\smaller\smaller$\scriptstyle\bullet$}}(P^{\text{\smaller\smaller$\scriptstyle\bullet$}},M^{\text{\smaller\smaller$\scriptstyle\bullet$}})\longrightarrow 0$ is exact. The total complex of any short exact sequence of complexes of abelian groups is exact. For a generalization, see Lemma 7.5 below. ∎
Lemma 2.6.
Let $\mathsf{B}$ be an abelian category with enough projective objects. Then any contraacyclic complex in $\mathsf{B}$ is acyclic.
Proof.
This is the dual assertion to [31, Lemma A.2] or a particular case of [29, Lemma B.7.3(b)]. It is not known whether the assertion of this lemma holds for exact categories in general; see [29, Remark B.7.4] for a discussion. ∎
A complex in an exact category $\mathsf{E}$ is said to be absolutely acyclic [27, Section 2.1], [28, Section 3.3], [29, Appendix A] if it belongs to the minimal thick subcategory of $\mathsf{Hot}(\mathsf{E})$ containing the total complexes of (termwise admissible) short exact sequences of complexes in $\mathsf{E}$. Equivalently, a complex in $\mathsf{E}$ is absolutely acyclic if and only if it belongs to the minimal full subcategory of $\mathsf{Com}(\mathsf{E})$ containing contractible complexes and closed under extensions and direct summands [41, Proposition 8.12].
@@
This is [27, Remark 2.1]. For additional information, see [34, Proposition 6.2]. ∎
Let us denote by $\mathsf{B}_{\mathsf{proj}}\subset\mathsf{B}$ the full subcategory of projective objects in an abelian/exact category $\mathsf{B}$.
@@
This is [40, Corollary 7.4]. A far-reaching generalization can be found in [41, Corollary 6.14]. See also Theorem 7.2 below. ∎
Our final lemma in this section is standard and easy, but helpful to keep in mind.
Lemma 2.9.
Let $\mathsf{H}$ be a triangulated category, $\mathsf{X}\subset\mathsf{H}$ be a thick subcategory, and $\mathsf{F}\subset\mathsf{H}$ be a full triangulated subcategory. Assume that for every object $H\in\mathsf{H}$ there exists an object $F\in\mathsf{F}$ together with a morphism $F\longrightarrow H$ whose cone belongs to $\mathsf{X}$. Then the inclusion functor $\mathsf{F}\longrightarrow\mathsf{H}$ induces a triangulated equivalence of the Verdier quotient categories
$\mathsf{F}/(\mathsf{X}\cap\mathsf{F})\simeq\mathsf{H}/\mathsf{X}.$
Proof.
This is [20, Proposition 10.2.7(ii)] or [28, Lemma 1.6(a)]. ∎